Find the solution to the system of equations by additonal method. X-2y=-7 -4x -1y=28 I would multiply 1 to get -1x =5y=-7 in the first equation right? Then I would add the second equation -4x-1y=28 to -1x=5y=-7 that gives me total of 5x 4y =35 right? So -7 is x, The next part is where I'm getting lost here. Would the answer be (-7,0)
Answer by drk(1908)

here are your two original equations:
(i) x - 2y = - 7
(ii) -4x -1y = 28
multiply (ii) by -2 to get
(i) x - 2y = - 7
(iii) + 8x + 2y = -56
add and solve for x.
(iv) 9x = -63
x = -7.
find y.
y = 0.
(-7,0)